Hafiz wanted to buy a brand-new car. He could only afford to pay RM 5000 for a down payment and decided to take a loan to finance the balance of the 139770
1
Expert's answer
2020-11-22T13:48:37-0500

Then let the loan be for 4 years at an annual rate of 10 %

find the annual payment amount

"S=\frac{L\times i}{1-\frac{1}{(1+i)^n}}=\frac{139770\times 0.1}{1-\frac{1}{(1+0.1)^4}}=\frac{13977}{0.316986545}=44093.35"

total loan amount:

"44093.35\times4=176373.4"